• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 14:47:22 on May29,2008

    Entry number 240307

    keyword=Cavities, part V. - Finished.

    Beam back at 1 uA. LNA installed. J. Musson decreases gain
    temporarily to adjust phases. (x,y) = +2, +2.

    Run 21050. T1 is 8.6 kHz / uA.
    Hence, for 50 nA we need 430 Hz.

    Next, we go for 50 nA. J. Musson maximizes cavity gains.
    MCC operators manage to put T1 to 530 Hz. It's pretty sensitive
    to the slits, so we'll run like this (no time for fine tuning).

    Run 21051, I_beam = 50 nA. MAT1H01H -2.33, MBC1C20V = 2.34,
    so (x,y) = (+2, +2).

    Run 21052, MAT1H01H = -2.33, MBC1C20V = +2.34, so (x,y) = (-2, +2).
    I'm confused about the readbacks (seems V changed, need to check
    offline).

    Musson finds the gains have saturated on Cav1. Change gain.
    Run again. Run 21053. (x,y) = -2, +2

    Return to +2,+2. Run 21054. This should be better than 21052
    because of the gains not saturated.

    Run at -2, -2. Run 21055. Cav1 Upstream Y position is saturating
    ocassionaly.

    Musson adjusts the gains again -- this time it should not saturate
    for any of the positions.
    Run 21056 at (-2,-2). T1 = 480 Hz. (135 Hz appears to be beam off)
    Run 21057 at (-2,+2). T1 = 500 Hz
    Run 21058 at (+2,+2). T1 = 650 Hz, probably angle changed.
    Run 21059 at (+2,-2). T1 = 450 Hz.

    According to the EPICS stripcharts there is plenty of signal and
    (online) a decent resolution at 50 nA. We can probably run
    at 10 nA, says Musson.

    14:46 Test finished.